\begin{thebibliography}{99}




\bibitem{BV}
O.A. Butkovsky, A.Yu. Veretennikov, 
On asymptotics for Vaserstein coupling of Markov chains, Stochastic Processes and their Applications, 123(9), 2013, 3518-3541.

\bibitem{Stroock-Diaconis}
P. Diaconis and D. Stroock, 
Ann. Appl. Probab., 1(1),  1991,  36-61.


\bibitem{Dobrushin}
R.L. Dobrushin, Central limit theorem for non-stationary Markov chains, I and II, Theory Prob. Appl., 1956, 1,  65-80, 329-383.  
%[English translation].

%\bibitem{Doeblin}
%W. Doeblin, Expos\'e de la th\'eorie des chaines simples constantes de Markov \`a un nombre fini d'\'etats. Math\'ematique de l'Union %Interbalkanique, 1938, 2, 77-105 \&  78-80.



\bibitem{Doob53}
J.L. Doob, Stochastic Processes, Wiley, 1953. 

\bibitem{Dynkin}
E.B. Dynkin, Markov processees: vol. 1, Springer, 2012. % (paperback). 


\bibitem{EthierKurtz} 
S.N. Ethier, T.G. Kurtz,  
Markov Processes: Characterization and Convergence, Wiley, 
%Series in Probability and Statistics, 
2005.

\bibitem{FW}
M.I. Freidlin, A.D. Wentzell, Random Perturbations of Dynamical Systems, Springer, 1984.

\bibitem{Gantmacher}
F.R. Gantmacher,  The Theory of Matrices. vol. 1, 2,   
AMS Chelsea Publ., 
Providence, Rhode Island, USA, 
2000.

\bibitem{Gnedenko}
B.V. Gnedenko, Theory of probability, 6th ed., Gordon and Breach Sci. Publ., Amsterdam, The Netherlands, et al., 1997.

\bibitem{Griffeath}
D.Z. Griffeath, A maximal coupling for Markov chains, 
Wahrscheinlichkeitstheorie verw. Gebiete,  1975, 31(2), 95-106.



%\bibitem{IbragimovLinnik}
%I.A. Ibragimov, Yu.V. Linnik, Independent and stationary sequences of random variables, Groningen: Wolters-Noordhoff Publishing, 1971.


\bibitem{Kalash}
V.V. Kalashnikov, Coupling Method, its Development and Applications. In the Russian translation of: E. Nummelin,  General Irreducible Markov Chains and Non-negative Operators (1984), Moscow, Mir, 1989.

%{\color{red}(something wrong: it is not a translation)}


\bibitem{KLS}
M.A. Krasnosel'skii,  E.A. Lifshits, A.V. Sobolev,  Positive Linear Systems: The method of positive operators, 
Berlin, Helderman Verlag, 1989. 


\bibitem{Lindvall}
T. Lindvall, Lectures on the Coupling Method, Dover, 2002.

\bibitem{Markov}
A.A. Markov, Extension of the limit theorems of probability theory to a sum of variables connected in a chain (1906, in Russian), reprinted in Appendix B of: R. Howard, Dynamic Probabilistic Systems, vol. 1: Markov Chains. John Wiley and Sons, 1971.

\bibitem{Nummelin}
E. Nummelin, 
General Irreducible Markov Chains,  %(Cambridge Tracts in Mathematics) (paperback), 
CUP, Cambridge, 2008. 


\bibitem{Seneta1}
E. Seneta, Non-negative Matrices and Markov Chains, 2nd ed., Springer, New York, 1981.

%\bibitem{Seneta2}
%E. Seneta, 
%Markov and the Birth of Chain Dependence Theory, International Statistical Review / Revue Internationale de Statistique, 1996, 64(3), %255--263.  
%DOI: 10.2307/1403785 


\bibitem{Thorisson}
H. Thorisson, Coupling, Stationarity, and Regeneration, Springer, New York, 2000.

%\bibitem{tutu}
%V.N. Tutubalin, Theory of probability and random processes, 1992, Moscow State University %Publ. House, Moscow  (in Russian). 
%%[CLT: Theorem 6.2.2]

\bibitem{Vaserstein}
L.N. Vaserstein, Markov Processes over Denumerable Products of Spaces, Describing Large Systems of Automata,  Problems of Information Transmission, 1969, 5(3), 47-52. 
 
\bibitem{Veretennikov17}
A.Yu. Veretennikov, Ergodic Markov processes and Poisson equations (lecture notes). In: Modern problems of stochastic analysis and statistics - Selected contributions in honor of Valentin Konakov (ed. by V. Panov), Springer, 2017, 457 - 511. 

\end{thebibliography}

\documentclass[12pt,english]{article}
\usepackage{geometry}
\usepackage{color}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}

\def\E{\hskip.15ex\mathrm{E}\hskip.10ex}
\def\P{\mathrm{P}}

\def\phi{\varphi}

\newif\ifabc
\abcfalse



\usepackage{babel}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}
\newtheorem{remark}{Remark}
\newtheorem{definition}
{Definition}
{Exercise}
{Assumption}
\newtheorem{corollary}
{Corollary}
\newtheorem{proposition}
{Proposition}

\newtheorem{example}
{Example}

\begin{document}
\global\long\def\E{\mathbb{E}}
\global\long\def\P{\mathbb{P}}
\global\long\def\N{\mathbb{N}}
\global\long\def\ind{\mathbb{I}}

\title{
{\normalsize\tt\hfill\jobname.tex}\\
On convergence rate for homogeneous Markov chains\\
}
\author{A.Yu. Veretennikov\footnote{ University of Leeds, UK; National Research University Higher School of Economics, and Institute for Information Transmission Problems, Moscow, Russia, email: a.veretennikov @ leeds.ac.uk. Supported by 
the grant RSF 17-11-01098}, 
M.A. Veretennikova\footnote{National Research University Higher School of Economics, Moscow, Russia, email: mveretennikova @ hse.ru. Supported by the  grant RSF 17-11-01098}
}

\maketitle
\begin{abstract}
This paper 
studies 
improved rates of convergence for ergodic homogeneous Markov chains. In comparison to the earlier papers the setting is also generalised to the case without a unique dominated measure.
Examples are provided where the new bound is compared with 
the classical Markov -- Dobrushin inequality and with the second eigenvalue of the transition matrix for finite state spaces. 
\end{abstract}


\section{Introduction} 
It is well-known that for a discrete irreducible acyclic (also called primitive) homogeneous Markov chain $(X_n, n\ge 0)$ with a finite state space $\cal S$ there is a unique stationary distribution $\mu$, and the distribution of the chain $\mu_n={\cal L}(X_n)$ admits an exponential bound uniformly with respect to the initial distribution $\mu_0$:
\begin{equation}\label{finiteMD}
\|\mu_n - \mu\|_{TV} \le 2 (1-\kappa)^n, \quad \forall n.
\end{equation}
This bound converges exponentially fast to zero if  
\begin{equation*}
\kappa = \min_{i,i'}\sum_{j\in {\cal S}} p_{ij}\wedge p_{i'j} >0, 
\end{equation*}
where $p_{ij}$ are the transition probabilities, see, e.g., \cite{Gnedenko}.  Here $\|\cdot\|_{TV}$ is the distance of total variation between measures. A similar bound exists for general Markov chains, too (cf., among many other sources, \cite{Veretennikov17}). Another well-known method of exponential estimates is related to eigenvalues of the transition probability matrix, see \cite[Chapter XIII, (96)]{Gantmacher}, \cite[Theorem 1.2]{Seneta1}. However, in the general situation it suits special cases where the eigenvectors (eigenfunctions) of the transition operator form a full basis in $L_2({\cal S})$. This approach works especially well  in the simplest case of $|{\cal S}| = 2$ (see, e.g., the Proposition \ref{pros2} below) and for symmetric cases which correspond to reversible Markov chains (see, e.g., \cite{Stroock-Diaconis}). 
In \cite{BV} and \cite{Veretennikov17} it was prompted that a certain markovian coupling (see \cite{Griffeath, Lindvall, Thorisson, Vaserstein} et al.) construction may in some cases provide a better convergence rate than that of (\ref{finiteMD}); note that this approach is applicable to non-reversible cases, too. Yet, neither paper provided examples \cite{BV,Veretennikov17}. Another small issue in both \cite{BV} and \cite{Veretennikov17} was a reasonable but still slightly restrictive an assumption of existence of a unique reference measure (in \cite{BV} the  Lebesgue one) for all transition kernels $Q(x, \cdot)= P(X_1 \in \cdot | X_0)|_{X_0=x}$. This assumption is dropped in the present paper. 

We emphasize that the eigenvalue method provides the best asymptotic convergence rate for finite transition matrices. However, the advantage of bound (\ref{finiteMD}) is that it is valid in general state spaces, see any texbook treating such a case, or, for example, \cite{Veretennikov17}, or chapter \ref{sec:erg_gen} below. Beside this generality, here some -- hopefully convincing -- examples are provided, which are inevitably finite (so far). 

These examples demonstrate that the new bound is effective -- unlike, e.g., the remarkable  Doeblin--Doob one \cite[Theorem V.5.6]{Doob53}. The only drawback of the latter is that the ``condition D'' from \cite{Doob53}, which is nowadays called the Doeblin-Doob condition, does not imply any computable bound on the constants in the resulting convergence rate, so this rate may actually be arbitrarily slow, albeit being exponential. The new approach  provides the same or a better bound in comparison to (\ref{finiteMD}).   
The authors are working on further examples for infinite state spaces.

The rest of the paper consists of two sections: 2 -- the presentation of the method from \cite{BV} and \cite{Veretennikov17} in general state spaces and under less restrictive assumptions (without a unique reference measure); 3 -- examples of both sorts, i.e., where the new bound is better and where it gives basically the same result as the classical MD inequality. 

\section{MD Ergodic Theorem and coupling, general case}\label{sec:erg_gen} 
We consider a homogeneous Markov process (MP) in discrete time $(X_n, \, n\ge 0)$ on a general state space $S$ with a topology and with its Borel sigma-algebra ${\cal B}(S)$.  
%
The following notations from the theory of Markov processes will be accepted (cf. \cite{Dynkin}): the index $x$ in $\mathbb E_x$ or $\mathbb P_x$ signifies the expectation or the probability measure related to the non-random initial state of the process $X_0$. This initial state may also be  random with some distribution $\mu$, in which case notations  $\mathbb E_\mu$ or $\mathbb P_\mu$ can be used. 

If  the state space $S$ is finite, then $|S|$ denotes the number of its elements and $\cal P$ stands for the transition matrix $\left(p_{ij}\right)_{1\le i,j \le |S|}$ of the process. Concerning the history of the MD bound see  \cite{Doob53, Dynkin, EthierKurtz, Seneta1}, et al.

For the most well-known inequality (\ref{exp_bd3}) of the Proposition \ref{thm_erg2} about the classical convergence rate bounds given below for the reader's convenience, it is assumed that the {\em Markov--Dobrushin constant} is positive:
\begin{equation}\label{MD}
 \kappa := \inf_{x,x'} \int \left(\frac{P_{x'}(1,dy)}{P_x(1,dy)}\wedge 1 \right)P_x(1,dy) > 0. 
\end{equation}
In the main result of this paper -- Theorem \ref{lastthm} in  the section -- this condition will be dropped. 
Note that here $\displaystyle \frac{P_{x'}(1,dy)}{P_x(1,dy)}$ is understood in the sense of the density of the absolute continuous components. For brevity we will be using a simplified notation $P_x(dz)$ for $P_x(1,dz)$. Note that integration of the kernel $P_x(1,dz)$ with any Borel function  is Borel measurable with respect to $x$ which is a standard requirement in Markov processes \cite{Dynkin}. An analogous measurability with respect to the pair $(x,x')$ will also be valid for an integral of a Borel function with the measure $\Lambda_{x,x'}$ defined by the formula
\begin{equation*}
\Lambda_{x,x'}(dz) := P_x(1,dz) + P_{x'}(1,dz), 
\end{equation*}  
due to linearity. 
Note that $\Lambda_{x,x'}(dz) = \Lambda_{x',x}(dz)$.
\begin{lemma}\label{newMD}
The following representation for the constant from (\ref{MD}) holds true: 
\begin{equation}\label{betterMD}
 \kappa = \inf_{x,x'} \int \left(\frac{P_{x'}(1,dy)}{\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{\Lambda_{x,x'}(dy)} \right)\Lambda_{x,x'} (dy) \;\; (>0). 
\end{equation}
\end{lemma}

\noindent
{\em Proof } -- straightforward.  
\ifabc
Let $\displaystyle f_{x,x'}(y) = \frac{P_x(1,dy)}{\Lambda_{x,x'} (dy)}(y)$. Then, 
\begin{eqnarray*}
&\displaystyle  \kappa = \inf_{x,x'} \int \left(\frac{P_{x'}(1,dy)}{P_x(1,dy)}\wedge \frac{P_{x}(1,dy)}{P_x(1,dy)} \right)P_x(1,dy) 
 \\\\
&\displaystyle  = \inf_{x,x'} \int \left(\frac{P_{x'}(1,dy)}{f_{x,x'}(y)\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{f_{x,x'}(y)\Lambda_{x,x'}(dy)} \right)f_{x,x'}(y)\Lambda_{x,x'} (dy) 
 \\\\
&\displaystyle = \inf_{x,x'} \int \left(\frac{P_{x'}(1,dy)}{\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{\Lambda_{x,x'}(dy)} \right)\Lambda_{x,x'} (dy), 
\end{eqnarray*}
as required. The Lemma \ref{newMD} is proved. 
\fi

\ifabc
\begin{remark}
Note that the right hand side in (\ref{newMD}), actually, does not depend on any particular reference measure $\Lambda_{x,x'}$ (even if it is not symmetric with respect to $x,x'$), i.e., for any other measure with respect to which both $P_{x'}(1,dy)$ and $P_{x}(1,dy)$ are absolutely continuous the formula (\ref{MD}) gives the same result. Indeed, it follows straightforward from the fact that if, say, $d\Lambda_{x,x'} <\!\!\!< d\tilde\Lambda_{x,x'}$ and $d\Lambda_{x,x'} = \phi_{x,x'} d\tilde\Lambda_{x,x'}$, then we get, 
\begin{eqnarray*}
&\displaystyle  \int \left(\frac{P_{x'}(1,dy)}{\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{\Lambda_{x,x'}(dy)} \right)\Lambda_{x,x'} (dy) 
 \\\\
&\displaystyle  = \int \left(\frac{P_{x'}(1,dy)}{\phi_{x,x'}\tilde\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{\phi_{x,x'}(y)\tilde\Lambda_{x,x'}(dy)} \right)\phi_{x,x'}(y) 1(\phi_{x,x'}(y)>0)\tilde\Lambda_{x,x'} (dy) 
 \\\\
&\displaystyle  = \int \left(\frac{P_{x'}(1,dy)}{\tilde\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{\tilde\Lambda_{x,x'}(dy)} \right)1(\phi_{x,x'}(y)>0)\tilde\Lambda_{x,x'} (dy).
\end{eqnarray*}
However, $P_{x'}(1,dy)<\!\!\!< \Lambda_{x,x'}(dy) = \phi_{x,x'}(y) \tilde \Lambda_{x,x'}(dy)$, so for any measurable $A$ we have $\int_A P_{x'}(1,dy) 1(\phi_{x,x'}(y)=0) = 0$ and the same for $P_x(1,dy)$, which means that, actually, 
\[
\int \left(\frac{P_{x'}(1,dy)}{\tilde\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{\tilde\Lambda_{x,x'}(dy)} \right)1(\phi_{x,x'}(y)>0)\tilde\Lambda_{x,x'} (dy) = \int \left(\frac{P_{x'}(1,dy)}{\tilde\Lambda_{x,x'}(dy)}\wedge \frac{P_{x}(1,dy)}{\tilde\Lambda_{x,x'}(dy)} \right)\tilde\Lambda_{x,x'} (dy).
\]
Respectively, if there are two reference measures $\Lambda_{x,x'}$ and, say, $\Lambda'_{x,x'}$, then we may take $\tilde\Lambda_{x,x'} = \Lambda_{x,x'} + \Lambda'_{x,x'}$, and the coefficients computed by using each of the two -- $\Lambda_{x,x'}$ and $\Lambda'_{x,x'}$ -- will be represented via $\tilde\Lambda_{x,x'}$ in the same way. 
\end{remark}

\fi

~

Here is the key notion in the following presentation: denote 
\[
\kappa(x,x') : = \int \left(\frac{P_{x'}(1,dy)}{P_x(1,dy)}\wedge 1 \right)P_x(1,dy).
\]

\ifabc
Also, let 
\[
\kappa^{(k)}(x,x') : = \int \left(\frac{P_{x'}(k,dy)}{P_x(k,dy)}\wedge 1 \right)P_x(k,dy).
\]
Clearly, for any $x,x'\in S$, 
\begin{equation*}
\kappa(x,x') \ge \kappa. 
\end{equation*}
\fi

\begin{lemma}\label{jj}
For any $x,x'\in S$,
\[
\kappa(x,x') = \kappa(x',x).
\]
\end{lemma}
\noindent
{\em Proof.} 
We have, 
 
\begin{eqnarray}
& \displaystyle \kappa(x',x) = \int \left(\frac{P_{x'}(1,dy)}{P_x(1,dy)}\wedge 1 \right)P_x(1,dy) 
= \int \left(\frac{P_{x'}(1,dy)}{P_x(1,dy)}\wedge 1 \right)\frac{P_x(1,dy)}{\Lambda_{x,x'}(dy)} \Lambda_{x,x'}(dy)
 \nonumber \\ \nonumber \\
& \displaystyle = \int \left(\frac{P_{x'}(1,dy)}{\Lambda_{x,x'}(dy)} \wedge \frac{P_x(1,dy)}{\Lambda_{x,x'}(dy)} \right) \Lambda_{x,x'}(dy). \label{newgamma}
\end{eqnarray}
The  Lemma \ref{jj} follows. 

\begin{definition}
We call Markov processes satisfying the condition (\ref{MD}) {\em Markov--Dobrushin's or {\bf MD}-processes}.
\end{definition}

In the scenario of finite homogeneous chains this condition (\ref{MD}) was introduced by Markov himself \cite{Markov}; later on, for non-homogeneous Markov processes its analogue was used by Dobrushin \cite{Dobrushin}. It was suggested by Seneta 
to call it the Markov--Dobrushin condition. Note that in all cases $\kappa \le 1$. The case $\kappa = 1$ corresponds to the i.i.d. sequence $(X_n, n\ge 0)$. In the opposite extreme situation where the transition kernels are singular for different $x$ and $x'$, we have $\kappa = 0$. 
The MD-condition (\ref{MD}) -- as well as (\ref{betterMD}) -- is most useful because it provides an {\bf effective} quantitative upper bound for the convergence rate of a Markov chain towards its (unique) invariant measure in total variation metric. The following theorem is classical: the bound (\ref{exp_bd3}) can be found in most textbooks on ergodic Markov chains. 

\begin{proposition}\label{thm_erg2}
Let the assumption (\ref{betterMD}) 
hold true. Then the process $(X_n)$ is ergodic, i.e., there exists  a limiting probability measure $\mu$, which is stationary and such that the uniform bound is satisfied for every $n$, 
\begin{equation}\label{exp_bd3}
 \sup_{x}\sup_{A\in {\cal B}(S)} |P_x(n,A) - \mu(A)| \le (1-\kappa)^{n}.
\end{equation}
\ifabc
Also, 
\begin{equation}\label{exp_bd4}
\sup_{x}\sup_{A\in {\cal B}(S)} |\mu_n(A) - \mu(A)| \le  (1-\kappa^{(k)})^{[n/k]}, 
\end{equation}
and 
\begin{equation}\label{exp_bd5}
\|\mu_n - \mu\|_{TV} \le 2 (1-\kappa^{(k)})^{[n/k]} (1-\kappa)^{n-k[n/k]}. 
\end{equation}
\fi
\end{proposition}
Equivalently, for any bounded Borel measurable function $g$ on $S$, 
\begin{equation}\label{exp_bd6}
\sup_x |\langle g,\mu^x_n\rangle - \langle g, \mu \rangle | \le |g|_B (1-\kappa^{})^{n},
\end{equation}
where $|g|_B = \sup_x |g(x)|$.
\ifabc
Note that ergodic theorem -- Proposition \ref{thm_erg2} -- implies the (weak and strong) law of large numbers (see, e.g., \cite[{\color{red}Theorem?}]{Seneta1}, or \cite{Veretennikov17}), 
\begin{equation}\label{exp_bd7}
\sup_x \mathbb P_x \left(\left|\frac1{n} \sum_{k=1}^n g(X_k) - \langle g, \mu \rangle\right| \ge \epsilon\right) \to 0, \quad n\to \infty,
\end{equation}
for any $\epsilon>0$, 
where  we only stated the weak form of the result.
\fi
\ifabc
Clearly, if the assumption (\ref{MD}) fails, the estimate (\ref{exp_bd3}) is still valid, but does not contain any information since the difference of two probabilities cannot exceed one in any case. Similarly, for  (\ref{exp_bd4}) and  (\ref{exp_bd5}) to make sense it is required that $\kappa^{(k)} > 0$, although, without this condition both inequalities are still valid. 
\fi

~

The following folklore lemma answers the following question: suppose we have two distributions, which are not singular and their ``common area'' equals some positive constant $\kappa$. Is it possible to realise these two distributions on one  probability space so that the two corresponding random variables {\em coincide} with probability equal exactly $\kappa$? 

\begin{lemma}[``Of two random variables'']\label{odvuh}
Let $\xi^{1}$ and $\xi^2$ be two random variables on their (different, without loss of generality) probability spaces $(\Omega^1, {\cal F}^1, \mathbb P^1)$ and $(\Omega^2, {\cal F}^2, \mathbb P^2)$ and with densities $p^1$ and $p^2$ with respect to some reference measure $\Lambda$, correspondingly.  Then, if 
\begin{equation*}
q := \int \left(p^1(x)\wedge p^2(x)\right) \Lambda(dx) > 0, 
\end{equation*}
then there exists one more probability space $(\Omega, {\cal F}, \mathbb P)$ and two random variables on it, $\eta^1, \eta^2$, such that 
\begin{equation*}
{\cal L}(\eta^j) ={\cal L}(\xi^j), \; j=1,2, \quad \& \quad \mathbb  P(\eta^1 = \eta^2) = q. 
\end{equation*}
\end{lemma}
For the proof of this simple well-known fact see, e.g., \cite{Veretennikov17}. Here $q$ plays the role of $\kappa$.

In this section it is explained how to apply the general  coupling method  to Markov chains in general state spaces $(S, {\cal B }(S))$. Various presentations of this method may be found in \cite{Kalash, Lindvall, Nummelin, Thorisson, Vaserstein},  et al. This section follows the lines from \cite{BV}, which, in turn, is based on \cite{Vaserstein}. Note that in \cite{BV} the state space was $\mathbb R^1$; however, in $\mathbb R^d$ all the formulae remain the same. 

Now, let us consider two versions of the same Markov process $(X^1_n), (X^2_n)$  with two initial distributions $\mu_0^1$ and  $\mu_0^2$ respectively (this does not exclude the case of non-random initial states). Denote 
\begin{equation*}
\kappa_0 = \kappa(\mu_0^1, \mu_0^2):=
\int \left(\frac{\mu_0^1(dy)}{\mu_0^2(dy)}\wedge 1 \right)\mu_0^2(dy).
\end{equation*}
It is clear that $0\le \kappa_0\le1$ similarly to $\kappa(u,v)$ for all $u,v$.  
If $X^1_0$ and $X^2_0$ have different distributions, then obviously $\kappa_0<1$. Otherwise, we  have
$X^1_n\stackrel{d}{=}X^2_n$ (equality in distribution) for all $n$, and the coupling can be made trivially by letting  $\widetilde X^1_n= \widetilde
X^2_n:=X^1_n$.

Let us introduce a new, vector-valued {\bf Markov process} $\left(\eta^1_n,\eta^2_n,\xi_n,\zeta_n\right)$.  If $\kappa_0=0$ then we set
\begin{equation*}
\eta^1_0:=X^1_0,\; \eta^2_0:=X^2_0,\; \xi_0:=0,\; \zeta_0:=1.
\end{equation*}
Otherwise, if $0<\kappa(0)<1$, then we apply the Lemma \ref{odvuh} with $\Lambda = \Lambda_{x,x'}$ conditional on $(X^1_0, X^2_0)=(x,x')$ to the random variables $X^1_0$ and $X^2_0$ so as to create the random variables 
$\eta^1_0$, $\eta^2_0$, $\xi_0$ and $\zeta_0$ (they correspond to $\eta^1, \eta^2, \xi$, and $\zeta$ in the Lemma); formally, this can be done if $\kappa_0=1$, too. Further, assuming that the random variables $\left(\eta^1_n,\eta^2_n,\xi_n,\zeta_n\right)$ have been determined for some $n$, let us show how to construct the vector $\left(\eta^1_{n+1},\eta^2_{n+1},\xi_{n+1},\zeta_{n+1}\right)$. For this aim, we define the transition probability density $\phi$ with respect to the measure $\Lambda_{x^1, x^2} \times \Lambda_{x^1, x^2} \times \Lambda_{x^1, x^2}\times (\delta_0 + \delta_1)$ for this vector-valued process as follows,
\begin{equation}\label{process_eta}
\phi(x,y):=\phi_1(x,y^1)\phi_2(x,y^2)\phi_3(x,y^3) \phi_4(x,y^4),
\end{equation}
where $x=(x^1,x^2,x^3,x^4)$, $y=(y^1,y^2,y^3,y^4)$, and if
 $0<\kappa(x^1,x^2)<1$, then
\begin{align}
&\displaystyle \phi_1(x,u):=\frac{p(x^1,u)-p(x^1,
u)\wedge p(x^2,u)}{1-\kappa(x^1,x^2)}, \quad
\phi_2(x,u):=\frac{p(x^2,u)-p(x^1,u)\wedge p(x^2,u)}{1-\kappa(x^1,x^2)},\label{phi_12}
 \\\nonumber\\
&\displaystyle \phi_3(x,u):=1(x^4=1)\frac{p(x^1,u)\wedge
 p(x^2,u)}{\kappa(x^1,x^2)}+1(x^4=0)p(x^3,u),\label{phi_3}
 \\\nonumber\\ 
&\displaystyle \phi_4(x,u):=1(x^4=1)\left(\delta_1(u)(1-\kappa(x^1,x^2))+ 
\delta_0(u)\kappa(x^1,x^2)\right) +1(x^4=0)\delta_0(u)\label{phi_4}, 
\end{align}
where $\delta_i(u)$ is the Kronecker symbol, $\delta_i(u) = 1(u=i)$, or, in other words, the delta measure concentrated at state $i$. The case $x^4=0$ signifies coupling which has already been realised at the previous step, and $u=0$ means successful coupling at the transition.  
In the degenerate cases, if $\kappa(x^1,x^2)=0$ (coupling at the transition is impossible), then instead of (\ref{phi_3}) we set,  e.g., 
$$
\phi_3(x,u):=1(x^4=1)p(x^3,u) + 1(x^4=0)p(x^3,u) = p(x^3,u), \qquad 
$$ 
and if $\kappa(x^1,x^2)=1$, then instead of (\ref{phi_12}) we may set 
$$
\phi_1(x,u)=\phi_2(x,u):= p(x^1,u). \qquad 
$$ 
The formula (\ref{phi_4}) which defines \(\phi_4(x,u)\) can be accepted in all cases.
\ifabc
In fact, in both degenerate cases $\kappa(x^1,x^2)=0$ or 
$\kappa(x^1,x^2)=1$, the functions $\phi_3(x,u)1(x^4=1)$ (or, respectively, $\phi_1(x,u)$ and $\phi_2(x,u)$) can be defined more or less arbitrarily, only so as to keep the property of conditional independence of the four random variables $\left(\eta^1_{n+1},\eta^2_{n+1},\xi_{n+1},\zeta_{n+1}\right)$ given $\left(\eta^1_n,\eta^2_n,\xi_n,\zeta_n\right)$. 
\fi

\begin{lemma}[\cite{BV}]\label{lemma:2}Let $X^1_0=x, X^2_0=x'$, 
and the random variables $\widetilde X^1_n$ and $\widetilde X^2_n$, for $n\in\mathbb{Z}_+$ be defined   by the following formulae:
\begin{align*}
\widetilde X^1_n:=\eta^1_n 1(\zeta_n=1)+\xi_n 1(\zeta_n=0), \quad 
\widetilde X^2_n:=\eta^2_n 1(\zeta_n=1)+\xi_n 1(\zeta_n=0).
\end{align*}
Then $\kappa_0 = \delta_{x,x'}$, and
\[
\widetilde X^1_n\stackrel{d}{=}X^1_n, \;\;\widetilde
 X^2_n\stackrel{d}{=}X^2_n, \quad \mbox{for all $n\ge 0$.}
\]
Moreover, 
\[
\widetilde X^1_n=\widetilde X^2_n, \quad \forall \; n\ge
n_0(\omega):=\inf\{k\ge0: \zeta_k=0\}, 
\] 
and
\begin{equation}\label{estimate}
\P_{x,x'}(\widetilde X^1_n\neq \widetilde
 X^2_n)\le(1-\kappa_0)\, \E_{x,x'}\prod_{i=0}^{n-1}
 (1-\kappa(\eta^1_i,\eta^2_i)).
\end{equation}
Moreover, $\left(\widetilde X^1_n\right)_{n\ge 0}$ and $\left(\widetilde X^2_n\right)_{n\ge 0}$ are both  homogeneous Markov processes, and 
\begin{equation*}
\left(\widetilde X^1_n\right)_{n\ge 0}\stackrel{d}{=}\left(X^1_n\right)_{n\ge 0}, \quad \& \quad 
\left(\widetilde X^2_n\right)_{n\ge 0}\stackrel{d}{=}\left(X^2_n\right)_{n\ge 0}.
\end{equation*}
\end{lemma}

\ifabc
Informally, the processes $\eta^1_n$ and $\eta^2_n$ represent $X^1_n$ and $X^2_n$, correspondingly,
under the condition that the coupling was not successful until time $n$, while the process $\xi_n$
represents both $X^1_n$ and $X^2_n$ if the coupling does occur no later than at time $n$. The process $\zeta_n$ represents the moment of coupling: the event  $\zeta_n=0$ is equivalent to the event that coupling occurs no later than at time $n$ while  $\zeta_n=1$ is the complementary event. 
As it follows from \eqref{process_eta} and \eqref{phi_4},
\begin{align*}
&\displaystyle \mathbb P(\zeta_{n+1}=0|\zeta_n=0)=1,
 \\ \\
&\displaystyle \mathbb P(\zeta_{n+1}=0|\zeta_n=1,\eta^1_n=x^1,\eta^2_n=x^2)=\kappa(x^1,x^2).
\end{align*}
Hence, if two processes were coupled at time $n$, then they remain coupled at time $n+1$, and if they were not
coupled, then the coupling occurs with the probability $\kappa(\eta^1_n,\eta^2_n)$.
At each time the probability of coupling at the next step is as large as possible, given the current states. 
\fi

\noindent
In \cite{BV} it was additionally assumed   in the Lemma \ref{lemma:2} that  there is a unique dominating measure for all transition kernels. This  assumption is not necessary and the proof runs without other  changes with $\Lambda_{x,x'}$ at each step instead of a unique dominating measure $\Lambda$. 

\begin{remark}\label{rem0}
If the initial values $X^1_0, X^2_0$ are distributed with laws $\mu^1_0, \mu_0^2$, correspondingly, then the following corollary  of the inequality (\ref{estimate}) holds, 
\begin{equation}\label{estimate2}
\P_{\mu^1_0,\mu^2_0}(\widetilde X^1_n\neq \widetilde
 X^2_n)\le(1-\kappa_0(\mu^1_0,\mu^2_0))\, \E_{\nu^1_0,\nu^2_0}\prod_{i=0}^{n-1}
 (1-\kappa(\eta^1_i,\eta^2_i)),
\end{equation}
where $\nu^j_0$ is the distribution of $\eta^j_0$, $j=1,2$. 

\end{remark}

~

Denote $\mbox{diag}(S^2): =  \{x\in S \times S: \, x^1=x^2\}$ and $\hat S^2 := S \times S \setminus \mbox{diag}(S^2)$. Consider the process $X_n := (\tilde X^1_n, \tilde X^2_n)$ on $S^2$ and its conditional version $\hat X_n := (\tilde X^1_n, \tilde X^2_n)|\tilde X^1_n \not = \tilde X^2_n$ on $\hat S^2$.
Note that $\hat X_n$ is also a Markov process. Indeed, its probabilities of states at time $n+1$ only depend on the state at time $n$, or, more precisely, given this state they do not depend on the past before $n$: in comparison to $X_n$ they are just recomputed so as to satisfy the normalisation condition.

~

Now using the function $\kappa(x)$, let us introduce an  operator $V$ acting on a (bounded, Borel measurable) function $h$ on the space $\hat S^2$ as follows: for $x=(x^1, x^2)\in  \hat S^2$,
\begin{equation}\label{V}
Vh(x) := (1-\kappa(x^1,x^2)) \mathbb E_{x^1,x^2}h(\hat X_1) \equiv \exp(\psi(x))\mathbb E_{x^1,x^2}h(\hat X_1),  
\end{equation}
where in the last expression $\psi(x):= \ln (1-\kappa(x^1,x^2))$ (assume $\ln 0 = -\infty$). From the estimate (\ref{estimate2}) it follows, 
\begin{equation}\label{l2}
\P_{\mu^1_0,\mu^2_0}(\widetilde X^1_n\neq \widetilde X^2_n)\le 
(1-\kappa_0(\mu^1_0,\mu^2_0)) \int V^n 1(x)\nu^1_0(dx^1)\nu^2_0(dx^2). 
\end{equation}
Note that by definition (\ref{V}), for the non-negative operator $V$ (which transforms any non-negative function into a non-negative one) its sup-norm $\|V\| = \|V\|_{B,B}:=\sup\limits_{|h|_B\le 1} |Vh|_B $ equals $\sup\limits_{x} V1(x)$, where 
$1=1(x)$ is the function on $\hat S^2$ identically equal to one. In our case 
$$
r(V) \le \|V\| = \sup\limits_{x} V1(x) = \sup_x \exp(\psi(x))\equiv \sup_x (1-\kappa(x)) = 1-\kappa, 
$$ 
due to the well-known inequality (see, for example, \cite[\S 8]{KLS}). 

Further, if the operator $V$ was compact (cf. \cite{KLS}), e.g., in the space $C(S)$ -- as is always true for finite $S$ -- then from the generalisation of the Perron--Frobenius Theorem (see, for example, \cite[\S 9, Theorem 9.2]{KLS}) it would follow (see, e.g., \cite[(7.4.10)]{FW}) that 
\begin{equation*}
\lim_{n\to\infty} \frac1n \, \ln V^n 1(x)  = \ln r(V). 
\end{equation*}
We also have, 
\(\sup_x V^n 1 (x) = \|V^n\|\) and  \(0\le  V^n 1 (x) \le  \|V^n\|\), for any \(x\in \hat S^2\). 
So from the Gelfand formula, 
\begin{equation}\label{l11}
\limsup_{n\to\infty} (V^n 1 (x))^{1/n} \le \lim_{n\to\infty} \|V^n\|^{1/n} =  r(V) \le \|V\|. 
\end{equation}
The latter inequality holds without any assumptions of compactness on $V$.
The assertions (\ref{l2}) and (\ref{l11}) together lead to the following result. 
\begin{theorem}\label{lastthm}
In all cases, 
\begin{equation*}
\limsup\limits_{n\to\infty} \frac1n \ln \| P_{x^1}(n,\cdot) - \mu(\cdot)\|_{TV} 
\le \limsup_{n\to\infty} \frac1n \, \ln \int V^n 1(x^1, x^2)\mu(dx^2)
\le \ln r(V).  
\end{equation*}
\end{theorem}
\begin{remark}\label{rem1}
\ifabc
If the process $(\hat X_n)$ is Markov ergodic {\color{red}(that is, satisfies the assumption (\ref{MD}) of the Proposition  
\ref{thm_erg2}?)}, then due to the law of large numbers there exists a limit 
\[
\lim_{n\to\infty} \frac1n \, \ln V^n 1(x), 
\] 
and this limit does not depend of $x$(??). In this case, 
\[
\lim_n \frac1n \, \ln V^n 1(x) = 
\sup_x \lim_n \frac1n \, \ln V^n 1(x) = r(V).  
\] 
\fi
If the analogue of the ergodicity condition (\ref{MD}) {\em is valid for the process $\hat X_n$,} 
then the strict inequality
\[
r(V) < \|V\| = 1-\kappa
\] 
holds iff $1-\kappa(\cdot)$ is {\em not a constant $\hat\mu$-a.s.,} where $\hat\mu$ is the unique invariant measure for $\hat X_n$.

\end{remark}

\ifabc

\begin{corollary}
Under the assumption
\begin{equation}\label{r1}
r(V)<1, 
\end{equation}
the rate of convergence in
\[
\|\mu_n - \mu\|_{TV} \to 0, \quad n\to\infty
\]
is exponential: 
for any $\epsilon>0$ and $n$ large enough ($n\ge N(x)$), 
\begin{equation}\label{newrate2}
\|P_x(n,\cdot) - \mu(\cdot)\|_{TV} \le (r(V)+ \epsilon)^n.  
\end{equation}
\end{corollary}

\noindent
To put it a little differently, if $r(V)<\|V\| = 1-\kappa$ and $\epsilon>0$ is chosen small enough so that $r(V)+ \epsilon < 1-\kappa$, then the bound (\ref{newrate2}) is strictly better than (\ref{exp_bd3}) for $n$ large enough. 

\begin{remark}
Let us emphasize that the bound in the Theorem \ref{lastthm} is asymptotic, for large $n$, unlike the strict bounds in the classical Ergodic Theorem and in the Diaconis--Stroock bound for reversible MC (see below).
\end{remark}

\begin{remark}
The condition (\ref{r1}) offers one possible (partial) answer to the question whether there is any intermediate situation in ``between'' Markov--Dobrushin's and Doeblin--Doob's with a bound like Doeblin--Doob's (see \cite{Doob53})
\begin{equation}\label{DD}
 \sup_{x}\sup_{A\in {\cal S}} |P_x(n,A) - \mu(A)| \le C\exp(-cn), \quad n\ge 0, 
\end{equation}
with some $C,c>0$, under the ``DD-condition'' which assumes that there exist a finite (sigma-additive) measure $\nu\ge 0$ and $\epsilon>0$, $s>0$ such that $\nu(A)\le \epsilon$ implies 
\[
\sup_x P_x(s, A) \le 1 - \epsilon. 
\]
The issue with the bound (\ref{DD}) is that the constants $C,c$ are not defined by the measure $\nu$ and the constant $\epsilon$. 

Beside the examples in the next section, note that if the MD condition $\kappa>0$ fails, it means $\kappa=0$, which just signifies that for at least one couple of states $i$ and $i'$ the kernels $Q_i(dy)$ and $Q_{i'}(dy)$ are singular, but it does not necessarily mean $r(V)=1$ since the process still may well be irreducible. So, indeed, the inequality $r(V)<1$ provides an intermediate condition more relaxed than MD and yet the one which allows an effective bound for the rate of convergence. 
\end{remark}


\fi


\section{Examples: Finite $S$}
\begin{proposition}\label{pros2}
Let us consider a $2 \times 2$ matrix $\cal{P}$ with elements $p_{1,1}=a, p_{1,2}=1-a, p_{2,1}=(1-b), p_{2,2}=b$, assuming $0<a,b<1$. 	Then $r=1-\kappa = |\lambda_{2}|$, where $\lambda_{2}$ is the smaller eigenvalue of the  matrix $\cal{P}$. 
\end{proposition}

{\em Proof} -- is straightforward and will be shown in the full version of the paper.  

~

In the examples below we compare the asymptotics provided by three methods: Markov--Dobrushin's, the one based on the second eigenvalue, and the one proposed in the Theorem \ref{lastthm} above. Although, in all examples the eigenvalues can be computed by hand, a code was written which can be applied to any Markov chain in a finite state; this code will be shown in the full version of the paper along with further multidimensional examples. 
\begin{example}[the three methods provide  similar rate bounds]
Consider the MP with the state space $S = \{1,2\}$ and a transition matrix
\[
{\cal P} = \left(
\begin{array}{l l}
0.65 & 0.35 \\
0.35 & 0.65
\end{array}
\right).
\]
We have, $\hat S^2 = \{(1,2), (2,1)\}$, and 
\[
V = (1-0.7) \times 
\left(
\begin{array}{l l}
0.65 - 0.35 & 0 \\
0 & 0.65 - 0.35
\end{array}
\right)\times \frac{1}{0.3} = 0.3 \times 
\left(
\begin{array}{l l}
1 & 0 \\
0 & 1
\end{array}
\right)\,.
\]
So we compute, 
\[
1-\kappa = 0.3   =  r(V), 
\]
as expected. We have, 
\[
\lambda_2 = 0.3 = r(V), 
\]
again as expected since here $\lambda_2 = 1-\kappa$.  
This Markov chain is reversible and we can compare our (asymptotic) bound with the more precise one from \cite[Proposition 3]{Stroock-Diaconis}:
\[
\|P_i(n,\cdot) - \mu\|_{TV} \le \left(\frac{1-\pi(i)}{2\pi(i)}\right)^{1/2} \lambda_2^n 
= \frac1{\sqrt{2}} \times 0.3^n. 
\]
Clearly, the asymptotics of the bounds is the same for all three approaches.


\end{example}

\begin{example}[The new approach is similar to the MD one; the eigenvalue one is better]
Consider the MP with the state space $S = \{1,2,3\}$ and transition matrix
\[
{\cal P} = \left(
\begin{array}{l l l}
0 & 0.3 & 0.7  \\
0.3 & 0.7 & 0 \\
0.7 & 0 & 0.3
\end{array}
\right)\,.
\]
Here the matrix is doubly stochastic, so the distrbution $\pi=(1/3, 1/3, 1/3)$ is invariant; hence, clearly, the process is reversible: $\pi_ip_{ij} = \pi_jp_{ji}$ for all $i,j$; $\kappa(.) = 0.3$, $1-\kappa = 0.7$. 

Simple computations show that $V=0.7  * \hat {\cal P}$ with some stochastic matrix $\hat {\cal P}$; the latter  is a transition matrix for the process $\hat X_n$ (see the previous section). Hence, its spectral radius $r(V)$ equals 0.7. That is, our bound asymptotically coincides with the classical one. 

The characteristic equation is 
\[
\lambda^{3}-\lambda^{2}-0.37\lambda = 0,
\]
and so the eigenvalues of $\cal P$ (without hat) are $\lambda_{1}=1, \lambda_{2,3}=\pm \sqrt{37}/10 \approxeq \pm 0.6082763$. Hence, we have $|\lambda_{2,3}| < r(V)$.

\end{example}

\begin{example}[The new approach is similar to the classical (MD); the eigenvalue one is better]
\[
{\cal P} = \left(
\begin{array}{l l l}
0 & 0.3 & 0.7  \\
0.7 & 0 & 0.3 \\
0.3 & 0.7 & 0
\end{array}
\right) \, \mbox{-- is doubly stochastic;} \; \pi = (1/3, 1/3,1/3) \; \mbox{is invariant.}; \; \kappa(.)\equiv0.3.
\]
Therefore, $1-\kappa=0.7$. The operator $V$ has a form 
$$
V = 0.7 \times \hat {\cal P}, \quad \mbox{with some stochastic transition matrix $\hat {\cal P}$}. 
$$
It is known that such a matrix has its spectral radius $r(V)=0.7$. Thus, in this example our new bound is similar to the classical one. 
\ifabc
One eigenvalue equals of $\cal P$ equals 1. The characteristic equation on eigenvalues reads, 
\[
-\lambda^3 +3 \lambda * 0.7 * 0.3 + 0.7^3 +0.3^3 = 0 \, \Leftrightarrow \,  \lambda_1=1, \lambda_{2,3} = - 0.25 \pm \sqrt{0.25 - 0.37} = - 0.25 \pm 0.12 i; 
\]
thus, 
\fi
The analysis shows the following:
\[
|\lambda_{2,3}| = \sqrt{0.25 + 0.12} = \sqrt{0.37}  \approxeq 0.6082763 < 0.7.
\]
Note that the matrix $\cal P$ (without hat) is not reversible ($\pi_i p_{ij} = \pi_j p_{ji}$ fails: e.g., $p_{12}/3 \not = p_{21}/3$), so, formally, the more precise bound from \cite[Proposition 3]{Stroock-Diaconis} is not applicable. 
\end{example}

\begin{example}[The new approach is better than the MD and similar to the eigenvalue one]
Consider the MP with the state space $S = \{1,2,3\}$ and transition matrix
\[
{\cal P} = \left(
\begin{array}{l l l}
0 & 0.3 & 0.7  \\
1.0 & 0 & 0 \\
0.8 & 0.1 & 0.1
\end{array}
\right)\,.
\]
Here $\hat S^2 = \{(1,2), (1,3), (2,3), (2,1), (3,1), (3.2)\} =: (I,II, III, IV, V, VI)$, 
\[
\kappa(1,2) =\kappa(2,1) = 0,\;  \kappa(1,3) =\kappa(3,1) = 0.2,\; \kappa(2,3) =\kappa(3,2) = 0.8;
\]
\[
1-\kappa(1,2)= 1-\kappa(2,1) = 1,\; 
1-\kappa(1,3) =1-\kappa(3,1) = 0.8,\; 1-\kappa(2,3) =1-\kappa(3,2) = 0.2.
\]
\ifabc
the non-normalised $\hat P$ is as follows, 
\[\hat 
{\cal P}_{nn} = \left(
\begin{array}{c c c c c c}
0 & 0 & 0 & 0.3 & 0.7 & 0 \\
0 & 0 & 0.03 & 0.24 & 0.56 & 0.07  \\
0.1 & 0.1 & 0  & 0 & 0 & 0 \\
0.3 & 0.7 & 0 & 0 & 0 & 0 \\
0.24 & 0.56 & 0.07 & 0 & 0 & 0.03 \\
0 & 0 & 0 & 0.1 & 0.1 & 0
\end{array}
\right)\,.
\]
the normalised $\hat P$ (divide all entries by the sum in each row):
\[\hat 
{\cal P}_{norm} = \left(
\begin{array}{c c c c c c}
0 & 0 & 0 & 0.3 & 0.7 & 0 \\
0 & 0 & 0.03/0.9 & 0.24/0.9 & 0.56/0.9 & 0.07/0.9  \\
0.5 & 0.5 & 0  & 0 & 0 & 0 \\
0.3 & 0.7 & 0 & 0 & 0 & 0 \\
0.24/0.9 & 0.56/0.9 & 0.07/0.9 & 0 & 0 & 0.03/0.9 \\
0 & 0 & 0 & 0.5 & 0.5 & 0
\end{array}
\right)\,.
\]
and finally (multiplying by $1-\kappa(x)$): 
$$
V=\left(
\begin{array}{c c c c c c}
0 & 0 & 0 & 0.3 & 0.7 & 0 \\
0 & 0 & 0.2666667 & 0.2133333 & 0.4977778 & 0.06222222  \\
0.1 & 0.1 & 0  & 0 & 0 & 0 \\
0.3 & 0.7 & 0 & 0 & 0 & 0 \\
0.2133333 & 0.4977778 & 0.06222222 & 0 & 0 & 0.02666667 \\
0 & 0 & 0 & 0.1 & 0.1 & 0
\end{array}
\right)\,.
$$
\fi 
The code -- to be published in the full version --  
provides the value $r(V)$: 
\[
1-\kappa=r(V) = 9/20+\sqrt{65}/20 \approx 0.85311289.
\]
The roots of the characteristic equation for the original transition probability matrix are $\lambda_{1}=1, \lambda_{2,3}=-9/20 \pm \sqrt{65}/20$, whilst the roots of the characteristic equation for the coupled process $\hat X_n$ are $\lambda_{c}=-9/20 - \sqrt{65}/20$, $\lambda_{c}=-9/20 + \sqrt{65}/20$, a repeated root $\lambda_{c}=0$, $\lambda_{c}=-\sqrt{65}/20 + 9/20$, $\lambda_{c}=\sqrt{65}/20 + 9/20$, where the last root has the highest value, i.e. $r=\sqrt{65}/20 + 9/20$ and so $|\lambda_{2}|=r=|-9/20 - \sqrt{65}/20| \approx 0.85311289$ indeed.
In this example the classical Markov -- Dobrushin method is useless since $1-\kappa=1$. The Remark \ref{rem1} is also applicable here.

\end{example}

\ifabc
\begin{example}
	Consider the MP with the state space ${\cal S} = \{1,2,3\}$ and transition matrix
	\[
	{\cal P} = \left(
	\begin{array}{l l l}
	0.1 & 0.1 & 0.8  \\
	0.3 & 0.7 & 0 \\
	0.6 & 0.4 & 0
	\end{array}
	\right)\,.
	\]
\ifabc
	
			Here $\hat S^2 = \{(1,2), (1,3), (2,1), (2,3), (3,1), (3,2)\} =: (I,II, III, IV, V, VI)$, 
	\[
	\kappa(1,2) =\kappa(2,1) = 0.2,\;  \kappa(1,3) =\kappa(3,1) = 0.5,\; \kappa(2,3) =\kappa(3,2) = 0.7;
	\]
	\[
	1-\kappa(1,2)= 1-\kappa(2,1) = 0.8,\; 
	1-\kappa(1,3) =1-\kappa(3,1) = 0.5,\; 1-\kappa(2,3) =1-\kappa(3,2) = 0.3;
	\]
	the non-normalised $\hat P$ is as follows, 
	\[\hat 
	{\cal P}_{nn} = \left(
	\begin{array}{c c c c c c}
	0.07 & 0 & 0.03 & 0 & 0.24 & 0.56 \\
	0.04 & 0 & 0.06 & 0 & 0.48 & 0.32  \\
	0.03 & 0.24 & 0.07  & 0.56 & 0 & 0 \\
	0.12 & 0 & 0.42 & 0 & 0 & 0 \\
	0.06 & 0.48 & 0.04 & 0.32 & 0 & 0 \\
	0.42 & 0 & 0.12 & 0 & 0 & 0
	\end{array}
	\right)\,.
	\]
	the normalised $\hat P$ (divide all entries by the sum in each row):
		\[\hat 
	{\cal P}_{nn} = \left(
	\begin{array}{c c c c c c}
	0.07/0.9 & 0 & 0.03/0.9 & 0 & 0.24/0.9 & 0.56/0.9 \\
	0.04/0.9 & 0 & 0.06/0.9 & 0 & 0.48/0.9 & 0.32/0.9  \\
	0.03/0.9 & 0.24/0.9 & 0.07/0.9  & 0.56/0.9 & 0 & 0 \\
	0.12/0.54 & 0 & 0.42/0.54 & 0 & 0 & 0 \\
	0.06/0.9 & 0.48/0.9 & 0.04/0.9 & 0.32/0.9 & 0 & 0 \\
	0.42/0.54 & 0 & 0.12/0.54 & 0 & 0 & 0
	\end{array}
	\right)\,.
	\]
	and finally (multiplying by $1-\kappa(x)$): 
	$$
	V=\left(
	\begin{array}{c c c c c c}
	0.06222222 & 0 & 0.02666667 & 0 &   0.2133333 & 0.4977778 \\
	002222222 & 0 & 0.03333333 & 0 & 0.2666667 & 0.1777778  \\
	0.02666667 & 0.2133333 & 0.06222222  & 0.49777778 & 0 & 0 \\
	0.06666667 & 0 & 0.23333333 & 0 & 0 & 0 \\
	0.03333333 & 0.2666667 & 0.02222222 & 0.1777778 & 0 & 0 \\
	0.23333333 & 0 & 0.06666667 & 0 & 0 & 0
	\end{array}
	\right)\,.
	$$
\fi	
	
	The code 
	provides the value $r(V)$: 
	\[
	1-\kappa  = 0.8  >  r(V) =  05152391.
	\]
	
\end{example}
\fi







\ifabc
\section{To add to the full file}

Let us consider a 2 by 2 matrix with elements $p_{1,1}=a, p_{1,2}=1-a, p_{2,1}=(1-b), p_{2,2}=b$ \\
 
~

{\bf Theorem?}\\
(or, at least a Proposition?)
 
~

Case 1) when $a, b \le 1/2$. 

Firstly, the second eigenvalue of this transition matrix has $|\lambda_{2}|=|1-(a+b)|$.

In this case $\kappa=a+b$, so $1-\kappa=1-(a+b)$. First we compute the transition matrix for the coupled process. Here, we will write I, II for states $(1,2), (2,1)$, respectively. We have that 
$$
p_{I, I}=p_{\{(1, 2), (2,1)\}}=\frac{a-\min(a,1-b)}{1-\kappa}*\frac{b-\min(b, 1-a)}{1-\kappa},
$$

and also 
$$
p_{I, II}=\frac{1-a-\min((1-a), b)}{1-\kappa}*\frac{(1-b)-\min((1-b), a)}{1-\kappa}.
$$

This boils down to $p_{I,I}=p_{II, II}=0$ and $p_{I, II}=p_{II, I}=1$. This is because for $a, b \le 1/2$ we have that $a-\min(a, 1-b)=a-a=0$, $1-\min(1-a, b)=1-(a+b)=1-\kappa$, $1-b-\min(1-b, a)=1-(a+b)=1-\kappa$,
$$
p_{I, I}=p_{(1,2), (1,2)}=\frac{a-\min(a, 1-b)}{1-\kappa}*\frac{b-\min(b, 1-a)}{1-\kappa}=0,
$$

$$
p_{II, II}=p_{(2, 1), (2,1)}=\frac{(b-\min(b, (1-a)))}{1-\kappa}*\kappa{a-\min(a, 1-b)}{1-\kappa}=b-b=0.
$$

Here the eigenvalues of $(1-\kappa)p$ thus also have the modulus of $|1-(a+b)|$, hence coinciding with the modulus of the second eigenvalue of the original transition matrix, as we claimed. 

Now, case (2a) is when $a \le 1/2 \le b$, $a \le 1-b$:

In this case we obtain $p_{I, I}=p_{II, II}=0, p_{I, II}=p_{II, I}=1$, because now $\min(a, 1-b)=a$ and $\min(1-a, b)=b$.

Case (2b): when $a \le 1/2 \le b$, $a > 1-b$:

In this case we obtain $p_{I, I}=p_{II, II}=1, p_{I, II}=p_{II, I}=0$, because now $\min(a, 1-b)=1-b$ and $min(b, 1-a)=1-a$.

Case 3: when $a, b \ge 1/2$:

In this case we obtain $p_{I, I}=p_{II, II}=1, p_{I, II}=p_{II, I}=0$, because now $\min(a, 1-b)=1-b$ and $\min(b, 1-a)=1-a$.
\fi

\ifabc
\section{Addendum (code in R) }
{\bf 1.} The code in {\bf R} for the Example 15.  
\begin{verbatim}
t = matrix(,  nrow=6, ncol=6)
t[1,1]=0
t[1,2]=0
t[1,3]=0
t[1,4]=0.3
t[1,5]=0.7
t[1,6]=0

t[2,1]=0
t[2,2]=0
t[2,3]=0.03/0.9
t[2,4]=0.24/0.9
t[2,5]=0.56/0.9
t[2,6]=0.07/0.9
  
t[3,1]=0.5
t[3,2]=0.5
t[3,3]=0
t[3,4]=0
t[3,5]=0
t[3,6]=0
  
  t[4,1]=0.3
  t[4,2]=0.7
  t[4,3]=0
  t[4,4]=0
  t[4,5]=0
  t[4,6]=0

  t[5,1]=0.24/0.9
  t[5,2]=0.56/0.9
  t[5,3]=0.07/0.9
  t[5,4]=0
  t[5,5]=0
  t[5,6]=0.03/0.9

  t[6,1]=0
  t[6,2]=0
  t[6,3]=0
  t[6,4]=0.5
  t[6,5]=0.5
  t[6,6]=0  


voq = matrix(,  nrow=6, ncol=1) \#This is the 1- kappa(x) vector for the example 15
>   
>   voq[1,1]=1
>   voq[2, 1]=0.8
>   voq[3,1]=0.2
>   voq[4,1]=1
>   voq[5,1]=0.8
>   voq[6,1]=0.2
>     
>   V6 = matrix(,  nrow=6, ncol=6) #This is V for example 15
>   
>     for(i in 1:6){
+       for(j in 1:6){
+         V6[i,j]=voq[i]*t[i,j]
+       }
+     }
>   V6
          [,1]      [,2]       [,3]      [,4]      [,5]       [,6]
[1,] 0.0000000 0.0000000 0.00000000 0.3000000 0.7000000 0.00000000
[2,] 0.0000000 0.0000000 0.02666667 0.2133333 0.4977778 0.06222222
[3,] 0.1000000 0.1000000 0.00000000 0.0000000 0.0000000 0.00000000
[4,] 0.3000000 0.7000000 0.00000000 0.0000000 0.0000000 0.00000000
[5,] 0.2133333 0.4977778 0.06222222 0.0000000 0.0000000 0.02666667
[6,] 0.0000000 0.0000000 0.00000000 0.1000000 0.1000000 0.00000000
>
> eigen(V6)
eigen() decomposition
$`values`
[1]  0.81406223+0.00000000i -0.79100980+0.00000000i -0.00814223+0.06558487i
[4] -0.00814223-0.06558487i -0.04592144+0.00000000i  0.03915346+0.00000000i

#Here we observe that the maximum eigenvalue of V is 0.81406223, and it is below 1.
\end{verbatim}

\begin{verbatim}
 
> t = matrix(,  nrow=6, ncol=6)
> t[1,1]=0.1*0.7/0.9
> t[1,2]=0
> t[1,3]=0.1*0.3/0.9
> t[1,4]=0
> t[1,5]=0.8*0.3/0.9
> t[1,6]=0.8*0.7/0.9
> 
> t[2,1]=0.1*0.4/0.9
> t[2,2]=0
> t[2,3]=0.1*0.6/0.9
> t[2,4]=0
> t[2,5]=0.8*0.6/0.9
> t[2,6]=0.8*0.4/0.9
> 
> t[3,1]=0.3*0.1/0.9
> t[3,2]=0.3*0.8/0.9
> t[3,3]=0.7*0.1/0.9
> t[3,4]=0.7*0.8/0.9
> t[3,5]=0
> t[3,6]=0
> 
> t[4,1]=0.3*0.4/0.54
> t[4,2]=0
> t[4,3]=0.7*0.6/0.54
> t[4,4]=0
> t[4,5]=0
> t[4,6]=0
> 
> t[5,1]=0.6*0.1/0.9
> t[5,2]=0.6*0.8/0.9
> t[5,3]=0.4*0.1/0.9
> t[5,4]=0.4*0.8/0.9
> t[5,5]=0
> t[5,6]=0
> 
> t[6,1]=0.6*0.7/0.54
> t[6,2]=0
> t[6,3]=0.4*0.3/0.54
> t[6,4]=0
> t[6,5]=0
> t[6,6]=0
> 
> voq = matrix(,  nrow=6, ncol=1)
> 
> voq[1,1]=0.8
> voq[2, 1]=0.5
> voq[3,1]=0.8
> voq[4,1]=0.3
> voq[5,1]=0.5
> voq[6,1]=0.3
> 
> V6 = matrix(,  nrow=6, ncol=6)
> 
> for(i in 1:6){
+   for(j in 1:6){
+     V6[i,j]=voq[i]*t[i,j]
+   }
+ }
> V6
[,1]      [,2]       [,3]      [,4]      [,5]      [,6]
[1,] 0.06222222 0.0000000 0.02666667 0.0000000 0.2133333 0.4977778
[2,] 0.02222222 0.0000000 0.03333333 0.0000000 0.2666667 0.1777778
[3,] 0.02666667 0.2133333 0.06222222 0.4977778 0.0000000 0.0000000
[4,] 0.06666667 0.0000000 0.23333333 0.0000000 0.0000000 0.0000000
[5,] 0.03333333 0.2666667 0.02222222 0.1777778 0.0000000 0.0000000
[6,] 0.23333333 0.0000000 0.06666667 0.0000000 0.0000000 0.0000000
> eigen(V6)
eigen() decomposition
$`values`
[1]  0.5152391 -0.3718102 -0.3279970  0.2881809  0.1683135 -0.1474818

$vectors
[,1]       [,2]        [,3]       [,4]       [,5]       [,6]
[1,] 0.5347483 -0.5067950  0.51722932  0.6059393  0.1643311  0.3031284
[2,] 0.3421982 -0.4376648  0.09310834  0.1001494 -0.6222570 -0.5392269
[3,] 0.5347483  0.5067950  0.51722932 -0.6059393  0.1643311 -0.3031284
[4,] 0.3113594 -0.2271746 -0.47307998 -0.3504392  0.2929020  0.3425603
[5,] 0.3421982  0.4376648  0.09310834 -0.1001494 -0.6222570  0.5392269
[6,] 0.3113594  0.2271746 -0.47307998  0.3504392  0.2929020 -0.3425603

> 
> t = matrix(,  nrow=6, ncol=6)
> t[1,1]=0.1*0.7
> t[1,2]=0
> t[1,3]=0.1*0.3
> t[1,4]=0
> t[1,5]=0.8*0.3
> t[1,6]=0.8*0.7
> 
> t[2,1]=0.1*0.4
> t[2,2]=0
> t[2,3]=0.1*0.6
> t[2,4]=0
> t[2,5]=0.8*0.6
> t[2,6]=0.8*0.4
> 
> t[3,1]=0.3*0.1
> t[3,2]=0.3*0.8
> t[3,3]=0.7*0.1
> t[3,4]=0.7*0.8
> t[3,5]=0
> t[3,6]=0
> 
> t[4,1]=0.3*0.4
> t[4,2]=0
> t[4,3]=0.7*0.6
> t[4,4]=0
> t[4,5]=0
> t[4,6]=0
> 
> t[5,1]=0.6*0.1
> t[5,2]=0.6*0.8
> t[5,3]=0.4*0.1
> t[5,4]=0.4*0.8
> t[5,5]=0
> t[5,6]=0
> 
> t[6,1]=0.6*0.7
> t[6,2]=0
> t[6,3]=0.4*0.3
> t[6,4]=0
> t[6,5]=0
> t[6,6]=0
> t
[,1] [,2] [,3] [,4] [,5] [,6]
[1,] 0.07 0.00 0.03 0.00 0.24 0.56
[2,] 0.04 0.00 0.06 0.00 0.48 0.32
[3,] 0.03 0.24 0.07 0.56 0.00 0.00
[4,] 0.12 0.00 0.42 0.00 0.00 0.00
[5,] 0.06 0.48 0.04 0.32 0.00 0.00
[6,] 0.42 0.00 0.12 0.00 0.00 0.00
> 

\end{verbatim}

\fi

\end{document}